OpenStudy (anonymous):

is sinx^2 equivilent to 1-cosx^2?

OpenStudy (apoorvk):

oh yes oh yes!!!

OpenStudy (apoorvk):

that is an identity. one o' the basic ones. ask any one in the order of phoenix, and they shall know!!! :p :p

sam (.sam.):

If (sinx)^2, then yes

OpenStudy (anonymous):

if this is what you meant... \[\sin^2x=1-\cos^2x\] the answer is yes... but... if this is what you meant.... \[\sin(x^2)=1-\cos(x^2)\] then the answer is no.